Which of the following is the equation of a line perpendicular to the line
y=-1/3x+1 , passing through the point (2,7)?
-3x-y=1
3x-y=-1
3x-y=1
-3x+y=-1

Answer by Alan3354(69443)

Which of the following is the equation of a line perpendicular to the line
y=-1/3x+1 , passing through the point (2,7)?
-3x-y=1
3x-y=-1
3x-y=1
-3x+y=-1
--------------
Check the slope. If it's perpendicular, the slope is +3.
Then, sub 2 for x and 7 for y, see if it's equal.